# Natural inflation in Palatini $F(R)$

### v1

@@ 1 Introduction
A promising approach is to embed inflation in the Palatini formulation of gravity (e.g. [28, 29, 30, 31] and refs. therein), in which the metric and connection are treated as independent variables. This framework modifies the way the scalar field interacts with gravity, potentially altering the predictions for inflation, often reducing the required field excursion and relaxing observational constraints from the CMB. Within this context, $F(R)$ theories provide a versatile extension [32, 33, 34, 35, 36, 37, 38], allowing the effective inflationary potential in the Einstein frame to be adjusted. In this work, we focus on models in which the potential remains bounded and positive, as in natural inflation, demonstrating that such scenarios can reconcile theory with observational data while preserving the desirable theoretical properties of the original model.
@@ 2 Palatini $F(R)$ theories
Finally, the number of e-folds can be expressed as follows
@@ 3 Natural inflation for $n\leq 2$
For the class of $F(R)$ models defined in Eq. (2.19), we obtain from Eq. (2.9) that
@@
where we remind the reader that $1<n<2$. These predictions correspond precisely to those of monomial inflation with power $\bar{n}=2(2-n)$ in the large-$N_{e}$ approximation. Therefore, when the “knee” is reached, the model effectively behaves like $\phi^{\bar{n}}$ inflation. As $\alpha$ increases, the predictions remain approximately stable around this configuration, but $\zeta_{N}$ (and consequently $\phi_{N}$ and $\chi_{N}$) slowly grows. Eventually, the shape of the potential in the inflationary region ceases to resemble a monomial form and becomes hilltop-like, explaining the subsequent turn towards lower values of $n_{s}$. To visualize all this in terms of the scalar field potential, we show in Fig. 4 the Einstein-frame potential $U(\chi)$ for $\alpha=0$ (blue), $\alpha=3.16\cdot 10^{6}$ (black, continuous), $\alpha=10^{7}$ (black, dashed), and $\alpha=10^{8}$ (black, dotted), with $M=500$ and $n=7/4$ (left), alongside a zoomed-in view of the same potential around the minimum (right). Notice that in Fig. 4 the potentials are normalized by
@@
with $\zeta_{\rm max}$ defined as the value of $\zeta$ satisfying
@@
Moreover, the Einstein-frame canonical field $\chi$ is normalized by $\chi_{\alpha}$, defined as
@@
so that it appears with a period of $2\pi$ in the plot. The subscript $\alpha$ emphasizes that $U_{\alpha}$ and $\chi_{\alpha}$ depend on $\alpha$. We indicate $\chi_{N}$ (stars) and $\chi_{\rm end}$ (dots) for $\alpha=0$ (lightest red), $3.16\cdot 10^{6}$ (lighter red), $10^{7}$ (red), and $10^{8}$ (darker red). The value $\alpha=3.16\cdot 10^{6}$ corresponds to the maximum of $n_{s}$, i.e., the “knee” in Fig. 3. For sufficiently small $\alpha$ [^2] It can be computed numerically that for $n=7/4$ and $M=500$, the higher-order contribution becomes relevant for $\alpha\gtrsim 3\cdot 10^{5}$ (see Figs. 2 and 3). the higher-order term is negligible, and the Einstein-frame potential coincides with the Jordan-frame potential. Once the higher-order term starts dominating, up to $\alpha\lesssim 3.16\cdot 10^{6}$, the Einstein-frame potential can be approximated by a monomial potential with $\bar{n}=2(2-n)$ in the inflationary region. For larger $\alpha$, this approximation breaks down, causing $n_{s}$ to decrease and producing the “tail” towards lower scalar spectral index values observed in panel (a) of Fig. 2 and 3.
Figure 4: The Einstein-frame potential $U(\chi)$ for $\alpha=0$ (blue), $\alpha=3.16\cdot 10^{6}$ (black, continuous), $\alpha=10^{7}$ (black, dashed), and $\alpha=10^{8}$ (black, dotted), with $M=500$ and $n=7/4$ (left), along with a zoomed-in view of the same potential (right). We indicate $\chi_{N}$ (stars) and $\chi_{\rm end}$ (dots) for $\alpha=0$ (lightest red), $3.16\cdot 10^{6}$ (lighter red), $10^{7}$ (red), and $10^{8}$ (darker red). The value $\alpha=3.16\cdot 10^{6}$ corresponds to the maximum of $n_{s}$, i.e., the “knee” in Fig. 3. The potentials are normalized by $U_{\alpha}$ (defined in Eq. (3.8)) and the Einstein-frame canonical field $\chi$ by $\chi_{\alpha}$ (defined in Eq. (3.10)), so that the potential appears with a period of $2\pi$ in the plot. Further details are given in the text.
In panel (d), the predictions for $\Lambda$ as a function of $\log_{10}(\alpha)$ are obtained by imposing $A_{s}\simeq 2.1\cdot 10^{-9}$. The behavior of $\Lambda$ directly reflects the “knee” structure seen in panel (a): it initially increases up to a maximum value
@@ 4 Natural inflation for $n>2$
We now consider the case with $n>2$. The CMB observables are still given by Eqs. (3.2) and (3.5). However, a crucial difference arises: the function $G(\zeta)$ is now positive only in the range
$0<\zeta<(\alpha(n-2))^{\frac{1}{1-n}}.$ (4.1)
At the same time, the function reaches its maximum at
@@
This represents a fundamental difference compared to the $n<2$ case, since we need to require $G(\zeta_{M})\geq 2\,\Lambda^{4}$ in order to ensure that the mapping between the Jordan-frame and Einstein-frame potentials is well defined. However, from Eqs. (4.2) and (4.3), we see that this condition cannot be satisfied for arbitrary values of $\alpha$, as $G(\zeta_{M})$ decreases with increasing $\alpha$. The maximum allowed value of $\alpha$ can therefore be computed as

### v2

@@ 1 Introduction
A promising approach is to embed inflation in the Palatini formulation of gravity (e.g. [28, 29, 30, 31] and refs. therein), in which the metric and connection are treated as independent variables. This framework modifies the way the scalar field interacts with gravity, potentially altering the predictions for inflation, often reducing the required field excursion and relaxing observational constraints from the CMB. Within this context, $F(R)$ theories provide a versatile extension (e.g. [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46] and refs. therein), allowing the effective inflationary potential in the Einstein frame to be adjusted. In this work, we focus on models in which the potential remains bounded and positive, as in natural inflation, demonstrating that such scenarios can reconcile theory with observational data while preserving the desirable theoretical properties of the original model.
@@ 2 Palatini $F(R)$ theories
Finally, the number of e-folds can be expressed as follows:
@@ 3 Natural inflation for $n\leq 2$
For the class of $F(R)$ models defined in Eq. (2.19), we obtain from Eq. (2.9) that:
@@
where we remind the reader that $1<n<2$. These predictions correspond precisely to those of monomial inflation with power $\bar{n}=2(2-n)$ in the large-$N_{e}$ approximation. Therefore, when the “knee” is reached, the model effectively behaves like $\phi^{\bar{n}}$ inflation. As $\alpha$ increases, the predictions remain approximately stable around this configuration, but $\zeta_{N}$ (and consequently $\phi_{N}$ and $\chi_{N}$) slowly grows. Eventually, the shape of the potential in the inflationary region ceases to resemble a monomial form and becomes hilltop-like, explaining the subsequent turn towards lower values of $n_{s}$. To visualize all this in terms of the scalar field potential, we show in Fig. 4 the Einstein-frame potential $U(\chi)$ for $\alpha=0$ (brown), $\alpha=3.16\cdot 10^{6}$ (black, continuous), $\alpha=10^{7}$ (black, dashed), and $\alpha=10^{8}$ (black, dotted), with $M=500$ and $n=7/4$ (left), alongside a zoomed-in view of the same potential around the minimum (right). Notice that in Fig. 4 the potentials are normalized by:
@@
with $\zeta_{\rm max}$ defined as the value of $\zeta$ satisfying:
@@
Moreover, the Einstein-frame canonical field $\chi$ is normalized by $\chi_{\alpha}$, defined as:
@@
so that it appears with a period of $2\pi$ in the plot. The subscript $\alpha$ emphasizes that $U_{\alpha}$ and $\chi_{\alpha}$ depend on $\alpha$. We indicate $\chi_{N}$ (stars) and $\chi_{\rm end}$ (dots) for $\alpha=0$ (blue), $3.16\cdot 10^{6}$ (yellow), $10^{7}$ (orange), and $10^{8}$ (red). The value $\alpha=3.16\cdot 10^{6}$ corresponds to the maximum of $n_{s}$, i.e., the “knee” in Fig. 3. For sufficiently small $\alpha$ [^2] It can be computed numerically that for $n=7/4$ and $M=500$, the higher-order contribution becomes relevant for $\alpha\gtrsim 3\cdot 10^{5}$ (see Figs. 2 and 3). the higher-order term is negligible, and the Einstein-frame potential coincides with the Jordan-frame potential. Once the higher-order term starts dominating, up to $\alpha\lesssim 3.16\cdot 10^{6}$, the Einstein-frame potential can be approximated by a monomial potential with $\bar{n}=2(2-n)$ in the inflationary region. For larger $\alpha$, this approximation breaks down, causing $n_{s}$ to decrease and producing the “tail” towards lower scalar spectral index values observed in panel (a) of Fig. 2 and 3.
Figure 4: The Einstein-frame potential $U(\chi)$ for $\alpha=0$ (brown), $\alpha=3.16\cdot 10^{6}$ (black, continuous), $\alpha=10^{7}$ (black, dashed), and $\alpha=10^{8}$ (black, dotted), with $M=500$ and $n=7/4$ (left), along with a zoomed-in view of the same potential (right). We indicate $\chi_{N}$ (stars) and $\chi_{\rm end}$ (dots) for $\alpha=0$ (blue), $3.16\cdot 10^{6}$ (yellow), $10^{7}$ (orange), and $10^{8}$ (red). The value $\alpha=3.16\cdot 10^{6}$ corresponds to the maximum of $n_{s}$, i.e., the “knee” in Fig. 3. The potentials are normalized by $U_{\alpha}$ (defined in Eq. (3.8)) and the Einstein-frame canonical field $\chi$ by $\chi_{\alpha}$ (defined in Eq. (3.10)), so that the potential appears with a period of $2\pi$ in the plot. Further details are given in the text.
In panel (d), the predictions for $\Lambda$ as a function of $\log_{10}(\alpha)$ are obtained by imposing $A_{s}\simeq 2.1\cdot 10^{-9}$. The behavior of $\Lambda$ directly reflects the “knee” structure seen in panel (a). it initially increases up to a maximum value:
@@ 4 Natural inflation for $n>2$
We now consider the case with $n>2$. The CMB observables are still given by Eqs. (3.2) and (3.5). However, a crucial difference arises: the function $G(\zeta)$ is now positive only in the range:
$0<\zeta<\zeta_{0},$ (4.1)
with $\zeta_{0}\equiv(\alpha(n-2))^{\frac{1}{1-n}}$. At the same time, the function reaches its maximum at:
@@
As it can be understood from Fig. 1, given the shape of $G(\zeta)$ we have two possibilities for mapping $\zeta$ to $\phi$. The first is to consider the range $0<\zeta<\zeta_{M}$, while the other is to consider $\zeta_{M}<\zeta<\zeta_{0}$. However, in the latter range it can be shown numerically that we always have $\epsilon(\zeta)\ll 1$ which doesn’t allow a graceful exit from inflation. For this reason, in what follows we only consider the $0<\zeta<\zeta_{M}$ range. Moreover, we need to require $G(\zeta_{M})\geq 2\,\Lambda^{4}$ in order to ensure that the mapping between the Jordan-frame and Einstein-frame potentials is well defined. This represents the main difference with the $n<2$ case. From Eqs. (4.2) and (4.3), we see that this condition cannot be satisfied for arbitrary values of $\alpha$, as $G(\zeta_{M})$ decreases with increasing $\alpha$. The maximum allowed value of $\alpha$ can therefore be computed as: